# Modelling Non-monotone Risk Aversion and Convex Compensation in Incomplete Markets

### v1

@@ A.5 Proof of Theorem 3
Similarly, for $\gamma_{n+1}^{-}=0$, we have for any $\xi_{t}$
$\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}=0,\quad g_{i}\left(\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}\right)=\infty,\quad\Phi^{\prime}\left(g_{i}\left(\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}\right)\right)=0,\quad\Phi\left(g_{i}\left(\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}\right)\right)=1,\quad i=0,1,2,$ (65)
@@
$X_{t}^{R}$ $\rightarrow e^{\left(-1+\frac{1}{\alpha_{n}}\right)\int_{t}^{T}\left(r_{s}+\frac{1}{2\alpha_{n}}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\,\right)\mathrm{d}s}\frac{1}{2}\left(\frac{\gamma_{n}}{y^{*}\xi_{t}}\right)^{\frac{1}{\alpha_{n}}}\left(1-0\right)\rightarrow\infty,$
$X_{t}^{\bar{R}}$ $\rightarrow e^{\left(-1-\frac{1}{\alpha_{n}}\right)\int_{t}^{T}\left(r_{s}-\frac{1}{2\alpha_{n}}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\,\right)\mathrm{d}s}\left(-\frac{1}{2}\right)\beta_{n}^{2}\left(\frac{\gamma_{n}}{y^{*}\xi_{t}}\right)^{-\frac{1}{\alpha_{n}}}\left(1-0\right)\rightarrow 0.$
@@
$X_{t}^{R}$ $\rightarrow e^{\left(-1+\frac{1}{\alpha_{0}}\right)\int_{t}^{T}\left(r_{s}+\frac{1}{2\alpha_{0}}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\,\right)\mathrm{d}s}\frac{1}{2}\left(\frac{\gamma_{0}}{y^{*}\xi_{t}}\right)^{\frac{1}{\alpha_{0}}}\left(1-0\right)\rightarrow 0,$
$X_{t}^{\bar{R}}$ $\rightarrow e^{\left(-1-\frac{1}{\alpha_{0}}\right)\int_{t}^{T}\left(r_{s}-\frac{1}{2\alpha_{0}}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\,\right)\mathrm{d}s}\left(-\frac{1}{2}\right)\beta_{0}^{2}\left(\frac{\gamma_{0}}{y^{*}\xi_{t}}\right)^{-\frac{1}{\alpha_{0}}}\left(1-0\right)\rightarrow-\infty.$
@@ A.6 Proof of Theorem 4
From the arguments in Appendix A.3, Problem (33) is equivalent to
@@ Appendix B Volatility estimation
There are a lot of models to study the properties of the volatility, including the stochastic and implied volatility models; see e.g., Berestycki, Busca and Florent (2002), Gatheral et al. (2010) and Cui, Kirkby and Nguyen (2018). Since this study does not focus on modeling the market, we do not adopt the complicated models. We describe three estimation methods of $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$ in the illustration on the performance of our optimal portfolio.
@@ B.2 Implied Volatility
Although the method of historical volatility is easy for implementation, it fails to capture some characteristics of Black–Scholes model. Another way to calculate $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$ is through the Black–Scholes formula for European option prices. In a multi-asset Black–Scholes model, let $P_{i,t}^{K}$ denote the price of a put option signed on the risky asset $S_{i,t}$ with a strike price $K$. According to Black and Scholes (1972), we have (with a slight abuse of notation)
@@
Hence, we can calculate the norm of the $ith$ row of $\boldsymbol{\sigma}_{t}$ and reallocate the weight to each component according to the correlation among the risky assets. However, in security markets, the value of $\|\boldsymbol{\sigma}_{i,t}\|_{2}^{2}$ varies corresponding to different $K$. Such phenomenon are referred to as the volatility smile or volatility smirk based on different scenarios. An interpretation of the volatility “smile” is the imperfection of financial markets. For example, put options on indexes usually incur the volatility “smirk”. Empirically, put options with higher strike prices imply lower volatility. Hence, the smirk can be viewed as the consequence of the bid–ask spread; see Peña, Rubio and Serna (2002). Since most traders long the indexes, they would like to hedge the risk by longing the put option on index-tracking ETFs. To hedge the extreme risks only, they tend to choose the put option with lower strike prices. This imbalance in demand results in an increase to put option prices with lower strike prices and a drop to the prices with higher strike prices. Hence, it is reasonable to estimate the real volatility by a weighted average of the implied volatilities calculated from put options with different levels of strike prices. As shown in Ederington and Guan (2002), the arithmetic average of implied volatility is good enough for prediction purposes. Since this study does not focus on econometrics, we simply choose the Euclidean norm of the $i$th asset’s implied volatility to be $\|\boldsymbol{\sigma}_{i,t}\|_{2}$. We use the historical return to get the correlation matrix of the risky assets, denoted by $\mathbf{C}$. Then we find the matrix $\mathbf{B}$ satisfying

### v2

@@ A.5 Proof of Theorem 3
Similarly, for $\gamma_{n+1}^{-}=0$, we have for any $\xi_{t}$,
$\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}=0,\quad g_{i}\left(\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}\right)=\infty,\quad\Phi^{\prime}\left(g_{i}\left(\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}\right)\right)=0,\quad\Phi\left(g_{i}\left(\frac{\gamma_{n+1}^{-}}{y^{*}\xi_{t}}\right)\right)=1,\quad i=0,1,2.$ (90)
@@
$X_{t}^{R}$ $\rightarrow\infty,\quad X_{t}^{\bar{R}}\rightarrow 0.$
@@
$X_{t}^{R}$ $\rightarrow 0,\quad X_{t}^{\bar{R}}\rightarrow-\infty.$
@@ A.6 Proof of Theorem 4
From the arguments in Section 4.2, Problem (67) is equivalent to the problem replaced with its concave envelope:
@@ Appendix B Volatility estimation
There are a lot of models to study the properties of the volatility, including the stochastic and implied volatility models; see e.g., Berestycki, Busca and Florent (2002), Gatheral et al. (2010), Dai, Tang and Yue (2016) and Cui, Kirkby and Nguyen (2018). Since this study does not focus on the market modeling, we do not adopt the complicated models. We describe three estimation methods of $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$ in the illustration on the performance of our optimal portfolio.
@@ B.2 Implied Volatility
Although the method of historical volatility is easy for implementation, it fails to capture some characteristics of the Black–Scholes model. Another way to calculate $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$ is through the Black–Scholes formula for European option prices. In a multi-asset Black–Scholes model, let $P_{i,t}^{K}$ denote the price of a put option signed on the risky asset $S_{i,t}$ with a strike price $K$. According to Black and Scholes (1972), we have (with a slight abuse of notation)
@@
Hence, we can calculate the norm of the $i$th row of $\boldsymbol{\sigma}_{t}$ and reallocate the weight to each component according to the correlation among the risky assets. However, in security markets, the value of $\|\boldsymbol{\sigma}_{i,t}\|_{2}^{2}$ varies corresponding to different $K$. Such phenomena are referred to as the volatility smile or volatility smirk based on different scenarios. An interpretation of the volatility “smile” is the imperfection of financial markets. For example, put options on indexes usually incur the volatility “smirk”. Empirically, put options with higher strike prices imply lower volatility. Hence, the smirk can be viewed as the consequence of the bid–ask spread; see Peña, Rubio and Serna (2002). Since most traders long the indexes, they would like to hedge the risk by longing the put option on index-tracking ETFs. To hedge the extreme risks only, they tend to choose the put option with lower strike prices. This imbalance in demand results in an increase to put option prices with lower strike prices and a drop to the prices with higher strike prices. Hence, it is reasonable to estimate the real volatility by a weighted average of the implied volatilities calculated from put options with different levels of strike prices. As shown in Ederington and Guan (2002), the arithmetic average of implied volatility is good enough for prediction purposes. Since this study does not focus on econometrics, we simply choose the $L^{2}$ norm of the $i$th asset’s implied volatility to be $\|\boldsymbol{\sigma}_{i,t}\|_{2}$. We use the historical return to get the correlation matrix of the risky assets, denoted by $\mathbf{C}$. Then we find the matrix $\mathbf{B}$ satisfying